# Rate-Distortion Theory for Deductive Sources under Closure Fidelity

### 20 Apr 2026 v2

@@ Abstract
We study lossy compression of a finite statement source generated in a fixed deductive environment. The source symbols are statements in a knowledge base endowed with a proof system, and reconstruction fidelity is measured by preservation of deductive closure rather than by symbolwise equality. This induces, once the proof system and canonical order are fixed, a decomposition of the source into an irredundant core and redundant stored consequences. Under a natural disjointness condition on zero-distortion reconstruction sets, we show that the minimum zero-distortion rate equals the source mass of the core times the entropy of the source conditioned on that core. For reconstruction alphabets contained in the deductive closure of the source knowledge base, we further prove that the full rate–distortion function depends only on the core, so redundant states are invisible to both rate and distortion. When the decoder is limited to a bounded inference-depth budget, we obtain an exact rate–depth–distortion characterization. Under an additional order-robustness assumption identifying the chosen core with the order-free essential set, this characterization interpolates between classical symbolwise compression and unconstrained deductive compression. These results formulate deductive compression as a structured source-coding problem and quantify how shared inference structure changes the fundamental limits of communication.
@@ II-A Deductive Source and Irredundant Core
We fix a finite ambient universe $\mathbb{S}_{O}$ of candidate statements together with an effective proof system $\mathsf{PS}$. For every finite $\Gamma\subseteq\mathbb{S}_{O}$ and every $s\in\mathbb{S}_{O}$, whether $\Gamma\vdash_{\mathrm{kb}}s$ is decidable. The induced deductive closure operator is
@@ II-B Closure Fidelity and Closure Distortion
with $0/0:=0$. In the sequel we use the single-letter distortion $d_{\text{Cn}}(s_{o},\hat{s}_{o}):=d_{\text{Cn}}(s_{o},\hat{s}_{o}\mid S_{O})$.
@@
For a deductive source $(S_{O},P_{O})$, a finite reconstruction alphabet $\hat{S}_{O}\subseteq\mathbb{S}_{O}$, and the single-letter distortion $d_{\text{Cn}}(\cdot,\cdot\mid S_{O})$, define
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \mathbb{E}[d_{\text{Cn}}(S_{o},\hat{S}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$ (2.2)
This is the rate–distortion function of the deductive source under closure fidelity.
@@ III Exact Zero-Distortion Rate
We adopt the convention $P_{A}H(\pi_{A}):=0$ when $P_{A}=0$.
@@
Using Theorem 3.1 and the classical identity $R(0;\,d_{H},\,P_{O})=H(P_{O})$ for zero Hamming distortion on a reconstruction alphabet containing $S_{O}$ [7, Sec. 10.5][8, Ch. 7], which is valid here because $S_{O}\subseteq\hat{S}_{O}$, completes the proof. ∎
@@ IV Core Decomposition of the Full Rate–Distortion Function
The key observation is that every redundant state contributes zero closure distortion against every reconstruction in $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, so both the distortion constraint and the mutual-information minimization reduce to the irredundant core. A lower bound follows by introducing the auxiliary variable $T(s_{o})=s_{o}$ on $A$ and $T(s_{o})=*$ on $J$, and then applying entropy concavity exactly as in the converse of Theorem 3.1. Achievability is obtained by extending an optimal test channel for the core sub-source to the redundant part using a common output law. The full proof is given in Appendix A-B. ∎
@@ V-A Finite-Step Deduction and Derivation Depth
    the sequence $T_{\mathsf{PS}}^{n}(B)$ stabilizes in finitely many steps.
@@ V-B Depth-Constrained Fidelity
Hence both $S_{O}\setminus\{s\}$ and $s$ belong to the latter set, so the whole source set $S_{O}$ is recovered within $\delta$ steps. Therefore $d_{\text{Cn}}^{\delta}(s,\hat{s}\mid S_{O})=0$. ∎
@@ V-C Exact Rate–Depth–Distortion Laws
Part (i) is Theorem 5.1 at $\delta=0$, where $A_{0}=S_{O}$, $P_{0}=1$, and $\pi_{0}=P_{O}$. Part (ii) follows from Proposition 5.1(iii) and Theorem 5.1. Indeed, for every $\delta\geq\mathsf{D_{d}}$, one has $A_{\delta}=A$, hence $P_{\delta}=P_{A}$ and $\pi_{\delta}=\pi_{A}$. Since $S_{O}\subseteq\hat{S}_{O}$, also $A_{\delta}=A\subseteq\hat{S}_{O}$, so the hypotheses of Theorem 5.1 are satisfied at each such $\delta$. Therefore
$R_{\mathrm{sem}}(0,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})=P_{A}H(\pi_{A}),\qquad\delta\geq\mathsf{D_{d}}.$
Part (iii) holds because the zero $\delta$-distortion feasible set enlarges with $\delta$, so the minimum mutual information cannot increase. ∎
@@
By Lemma 5.1, every symbol outside $A_{\delta}$ is distortion-free under $d_{\text{Cn}}^{\delta}$, so both the distortion constraint and the mutual-information minimization reduce to the $\delta$-core. The lower bound follows by introducing $T_{\delta}(s_{o})=s_{o}$ on $A_{\delta}$ and $T_{\delta}(s_{o})=*$ on its complement, exactly as in the proof of Theorem 4.1. Achievability is obtained by extending an optimal test channel for the $\delta$-core sub-source to the $\delta$-redundant part using a common output law. At $\delta=0$, one has $A_{0}=S_{O}$ and $d_{\text{Cn}}^{0}=d_{H}$, so the formula reduces to the classical Hamming rate–distortion function [25, 7, 8]. Full details are given in Appendix A-E. ∎
@@ VI Consequences: Separation, Converse Bounds, and One Heterogeneous Extension
The preceding sections are source-coding results. We now record three consequences. First, because the relevant distortions are single-letter and bounded, the classical source–channel separation theorem applies verbatim [25, 7, 8, 22]. Second, the core/redundant decomposition yields sharper converse bounds than symbolwise fidelity. Third, the same zero-distortion mechanism extends to a restricted reconstruction alphabet, giving a compact heterogeneous-receiver extension.
@@ VI-A Separation and Operational Blocklength
Each distortion family considered here is a bounded single-letter distortion on a finite alphabet. Therefore the standard source–channel separation theorem for discrete memoryless sources and channels applies; see, e.g., [7, Chs. 7 and 10], [8, Ch. 3], [22, Chs. 7 and 8]. For $d_{\text{Cn}}$, the relevant source-coding function is $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$; for $d_{\text{Cn}}^{\delta}$, it is $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$. The zero-distortion thresholds follow by substituting the exact expressions from Theorems 3.1 and 5.1. Full details are given in Appendix A-F. ∎
@@
Part (i) is the standard Fano converse [7, Sec. 2.10][22, Sec. 1.9].
For (ii), let $M_{A}$ be uniformly distributed on $A$ and consider the induced $|A|$-message subproblem. Since $P_{e,\text{Cn}}^{(n)}\leq\epsilon$ in the max-error sense, the average closure error probability for $M_{A}$ is also at most $\epsilon$. Under Assumption 3.1, the zero-distortion reconstruction sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so a decoder output that is closure-correct for one core symbol cannot be closure-correct for another. Thus closure-correct decoding of core symbols induces an ordinary $|A|$-ary message recovery problem, and Fano’s inequality yields (6.10). ∎
@@
Part (i) is Shannon channel coding for a message set of size $|S_{O}|$; see, e.g., [26, 7, 8].
For (ii), code only the core $A$. Choose an $(n,|A|)$ reliable channel code for $A$. If the transmitted symbol is $a\in A$, decode it as $a$. If the transmitted symbol is $j\in J$, encode it using the codeword of any fixed $a_{0}\in A$, and decode to whatever core symbol the channel decoder outputs. Closure correctness for $a\in A$ follows from correct decoding, and closure correctness for $j\in J$ follows from Proposition 2.2. ∎
@@ VI-B Closure-Adapted Converse Bounds
Because $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, every redundant input is automatically closure-correct by Proposition 2.2, so closure errors can arise only on the core $A$. Conditioning on the event $B:=\mathbf{1}[S_{o}\in A]$, one reduces the problem to a $|A|$-ary source with law $\pi_{A}$. Under Assumption 3.1, the sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so the decoder output induces an ordinary decision rule on $A\cup\{\mathsf{e}\}$. Applying the classical Fano inequality to the conditioned core source and then multiplying by $P_{A}$ yields the bound. Full details are given in Appendix A-G. ∎
@@ VI-C One Heterogeneous Extension
For the converse, any zero-distortion test channel supported on $V$ must map each $a\in A$ into $R_{\text{Cn}}^{(V)}(a)$, and these sets are pairwise disjoint by (H2). Thus the converse proof of Theorem 3.1 applies verbatim after replacing $R_{\text{Cn}}(a)$ by $R_{\text{Cn}}^{(V)}(a)$. Full details are given in Appendix A-H. ∎
@@ Appendix A Proofs of Sections III–VI
Since $T$ is a deterministic function of $S_{o}$, the data processing inequality gives
@@
Since the supports of $\{P_{a}\}_{a\in A}$ are pairwise disjoint, the disjoint-support entropy identity yields
@@
The output marginal is $P_{A}\bar{\pi}+P_{J}Q$, and concavity of entropy gives
@@
Then $T$ is a deterministic function of $S_{o}$, so
@@
Step 3: achievability by homogenizing the redundant part. Let $P_{\hat{S}\mid A}^{*}$ achieve $R^{(A)}(D/P_{A};\,d_{\text{Cn}},\,\pi_{A})$, and let
@@
Since $T_{\delta}$ is a deterministic function of $S_{o}$,
@@
Since the supports of $\{P_{a}\}_{a\in A_{\delta}}$ are pairwise disjoint, the disjoint-support entropy identity gives
@@
A-D Proof of Corollary 5.1
$R_{\mathrm{sem}}(0,0;\,d_{\text{Cn}}^{0},\,P_{O})=H(P_{O}).$
For part (iii), note that if $\delta_{1}\leq\delta_{2}$, then
So $\delta\mapsto R_{\mathrm{sem}}(0,\delta)$ is non-increasing. ∎
A-E Proof of Theorem 5.2
@@
As in the proof of Theorem 4.1,
$I(S_{o};\hat{S}_{o})\geq I(T_{\delta};\hat{S}_{o})\geq P_{\delta}\,I(\mathsf{A}_{\delta,o};\hat{\mathsf{S}}_{o}),$
@@
For achievability, let $P_{\hat{S}\mid A_{\delta}}^{*}$ achieve the right-hand side, and extend it to all of $S_{O}$ by assigning to every $j\notin A_{\delta}$ the output marginal induced by $P_{\hat{S}\mid A_{\delta}}^{*}$. This keeps the distortion unchanged because the symbols outside $A_{\delta}$ are $\delta$-redundant, and the same entropy calculation as in Appendix A-B shows that the resulting mutual information is exactly
@@
A-F Proof of Theorem 6.1
All distortion functions used in the theorem are bounded single-letter distortions on finite alphabets. Therefore the standard source–channel separation theorem applies.
For closure distortion $d_{\text{Cn}}$, the source-coding side is governed by $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$, and the channel-coding side is governed by the Shannon capacity $C(W)$. Hence a distortion level $D$ is achievable whenever
@@
Let $B:=\mathbf{1}[S_{o}\in A]$. Since $B$ is a deterministic function of $S_{o}$,
@@
Because the sets $R_{\text{Cn}}(a)$ are pairwise disjoint, $g$ is well-defined. By data processing,
@@
Let $p_{A}:=\epsilon_{\text{Cn}}/P_{A}$. Applying the classical Fano inequality [7, Sec. 2.10][22, Sec. 1.9] to the $|A|$-ary source conditioned on $B=1$ gives

### 25 Apr 2026 v3

@@ Abstract
We study lossy compression of a finite statement source generated in a fixed deductive environment. The source symbols are statements in a knowledge base endowed with a proof system, and reconstruction fidelity is measured by preservation of deductive closure rather than by symbolwise equality. This induces, once the proof system and canonical order are fixed, a decomposition of the source into an irredundant core and redundant stored consequences. Under a natural disjointness condition on zero-distortion reconstruction sets, we show that the minimum zero-distortion rate equals the source mass of the core times the entropy of the source conditioned on that core. For reconstruction alphabets contained in the deductive closure of the source knowledge base, we further prove that the full rate–distortion function depends only on the core, so redundant states are invisible to both rate and distortion. When the decoder is limited to a bounded inference-depth budget (a bounded number of iterations of the immediate-consequence operator), we obtain an exact rate–depth–distortion characterization. Under an additional order-robustness assumption identifying the chosen core with the order-free essential set, this characterization interpolates between classical symbolwise compression and unconstrained deductive compression. These results formulate deductive compression as a structured source-coding problem and quantify how shared inference structure changes the fundamental limits of communication.
@@ II-A Deductive Source and Irredundant Core
We fix a finite ambient universe $\mathbb{S}_{O}$ of candidate statements together with an effective proof system $\mathsf{PS}$. Hence $\text{Cn}(\Gamma)\subseteq\mathbb{S}_{O}$ is finite for every finite $\Gamma$. For every finite $\Gamma\subseteq\mathbb{S}_{O}$ and every $s\in\mathbb{S}_{O}$, whether $\Gamma\vdash_{\mathrm{kb}}s$ is decidable. The induced deductive closure operator is
@@ II-B Closure Fidelity and Closure Distortion
with $0/0:=0$, i.e., zero distortion when both closure sets are empty. In the sequel we use the single-letter distortion $d_{\text{Cn}}(s_{o},\hat{s}_{o}):=d_{\text{Cn}}(s_{o},\hat{s}_{o}\mid S_{O})$.
@@
Let $(S_{O},P_{O})$ be a deductive source. Let $\mathsf{S}_{o}\sim P_{O}$ and let $\hat{\mathsf{S}}_{o}\in\hat{S}_{O}$ be generated via a test channel $P_{\hat{S}\mid S}$ as $\hat{\mathsf{S}}_{o}\sim P_{\hat{S}\mid S}(\cdot\mid\mathsf{S}_{o})$. With the single-letter distortion $d_{\text{Cn}}(\cdot,\cdot\mid S_{O})$, define
$R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O}):=\min_{\begin{subarray}{c}P_{\hat{S}\mid S}:\\ \mathbb{E}[d_{\text{Cn}}(\mathsf{S}_{o},\hat{\mathsf{S}}_{o})]\leq D\end{subarray}}I(\mathsf{S}_{o};\hat{\mathsf{S}}_{o}).$ (2.2)
@@ III Exact Zero-Distortion Rate
We write $h_{b}(p):=-p\log p-(1-p)\log(1-p)$ for the binary entropy function, and adopt the convention $P_{A}H(\pi_{A}):=0$ when $P_{A}=0$.
@@
Using Theorem 3.1 and the classical identity $R(0;\,d_{H},\,P_{O})=H(P_{O})$ for zero Hamming distortion on a reconstruction alphabet containing $S_{O}$ [7, Ch. 5][8, Ch. 7], which is valid here because $S_{O}\subseteq\hat{S}_{O}$, completes the proof. ∎
@@ IV Core Decomposition of the Full Rate–Distortion Function
By Proposition 2.2, every redundant state contributes zero closure distortion against every reconstruction in $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, so both the distortion constraint and the mutual-information minimization reduce to the irredundant core. A lower bound follows by introducing the auxiliary variable $T(s_{o})=s_{o}$ on $A$ and $T(s_{o})=*$ on $J$, and then applying entropy concavity exactly as in the converse of Theorem 3.1. Achievability is obtained by extending an optimal test channel for the core sub-source to the redundant part using a common output law. The full proof is given in Appendix A-B. ∎
@@ V-A Finite-Step Deduction and Derivation Depth
    the sequence $T_{\mathsf{PS}}^{n}(B)$ stabilizes in finitely many iterations.
@@ V-B Depth-Constrained Fidelity
Hence both $S_{O}\setminus\{s\}$ and $s$ belong to the latter set, so the whole source set $S_{O}$ is recovered within $\delta$ iterations. Therefore $d_{\text{Cn}}^{\delta}(s,\hat{s}\mid S_{O})=0$. ∎
@@ V-C Exact Rate–Depth–Distortion Laws
$R_{\mathrm{sem}}(0,0;\,d_{\text{Cn}}^{0},\,P_{O})=H(P_{O}).$
For part (iii), note that if $\delta_{1}\leq\delta_{2}$, then
So $\delta\mapsto R_{\mathrm{sem}}(0,\delta)$ is non-increasing. ∎
@@
By Lemma 5.1, every symbol outside $A_{\delta}$ is distortion-free under $d_{\text{Cn}}^{\delta}$, so both the distortion constraint and the mutual-information minimization reduce to the $\delta$-core. The lower bound follows by introducing $T_{\delta}(s_{o})=s_{o}$ on $A_{\delta}$ and $T_{\delta}(s_{o})=*$ on its complement, exactly as in the proof of Theorem 4.1. Achievability is obtained by extending an optimal test channel for the $\delta$-core sub-source to the $\delta$-redundant part using a common output law. At $\delta=0$, one has $A_{0}=S_{O}$ and $d_{\text{Cn}}^{0}=d_{H}$, so the formula reduces to the classical Hamming rate–distortion function [24, 7, 8]. Full details are given in Appendix A-D. ∎
@@ VI Consequences: Separation, Converse Bounds, and One Heterogeneous Extension
The preceding sections are source-coding results. We now record three consequences. First, because the relevant distortions are single-letter and bounded, the classical source–channel separation theorem applies verbatim [24, 7, 8]. Second, the core/redundant decomposition yields sharper converse bounds than symbolwise fidelity. Third, the same zero-distortion mechanism extends to a restricted reconstruction alphabet, giving a compact heterogeneous-receiver extension.
@@ VI-A Separation and Operational Blocklength
Each distortion family considered here is a bounded single-letter distortion on a finite alphabet. Therefore the standard source–channel separation theorem for discrete memoryless sources and channels applies. For $d_{\text{Cn}}$, the relevant source-coding function is $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$; for $d_{\text{Cn}}^{\delta}$, it is $R_{\mathrm{sem}}(D,\delta;\,d_{\text{Cn}}^{\delta},\,P_{O})$. The zero-distortion thresholds follow by substituting the exact expressions from Theorems 3.1 and 5.1. Full details are given in Appendix A-E. ∎
@@
Part (i) is the standard Fano converse [7, Ch. 2][8, Ch. 3].
For (ii), let $M_{A}$ be uniformly distributed on $A$ and consider the induced $|A|$-message subproblem. Since $P_{e,\text{Cn}}^{(n)}\leq\epsilon$ in the max-error sense, the average closure error probability for $M_{A}$ is also at most $\epsilon$ by definitions of max-error and average error [7, Ch. 7][8, Ch. 6]. Under Assumption 3.1, the zero-distortion reconstruction sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so a decoder output that is closure-correct for one core symbol cannot be closure-correct for another. Thus closure-correct decoding of core symbols induces an ordinary $|A|$-ary message recovery problem [7, Ch. 7][8, Ch. 11], and Fano’s inequality yields (6.10). ∎
@@
Part (i) is Shannon channel coding for a message set of size $|S_{O}|$; see, e.g., [7, Ch. 7][8, Ch. 6].
For (ii), code only the core $A$. Choose an $(n,|A|)$ reliable channel code for $A$ [7, Ch. 7][8, Ch. 6]. If the transmitted symbol is $a\in A$, decode it as $a$. If the transmitted symbol is $j\in J$, encode it using the codeword of any fixed $a_{0}\in A$, and decode to whatever core symbol the channel decoder outputs. Closure correctness for $a\in A$ follows from correct decoding, and closure correctness for $j\in J$ follows from Proposition 2.2. ∎
@@ VI-B Closure-Adapted Converse Bounds
Because $\hat{S}_{O}\subseteq\text{Cn}(S_{O})$, every redundant input is automatically closure-correct by Proposition 2.2, so closure errors can arise only on the core $A$. Conditioning on the event $B:=\mathbf{1}[S_{o}\in A]$, one reduces the problem to a $|A|$-ary source with law $\pi_{A}$. Under Assumption 3.1, the sets $\{R_{\text{Cn}}(a):a\in A\}$ are pairwise disjoint, so the decoder output induces an ordinary decision rule on $A\cup\{\mathsf{e}\}$. Applying the classical Fano inequality to the conditioned core source and then multiplying by $P_{A}$ yields the bound. Full details are given in Appendix A-F. ∎
@@ VI-C One Heterogeneous Extension
For the converse, any zero-distortion test channel supported on $V$ must map each $a\in A$ into $R_{\text{Cn}}^{(V)}(a)$, and these sets are pairwise disjoint by (H2). Thus the converse proof of Theorem 3.1 applies verbatim after replacing $R_{\text{Cn}}(a)$ by $R_{\text{Cn}}^{(V)}(a)$. Full details are given in Appendix A-G. ∎
@@ Appendix A Proofs of Sections III–VI
Since $T$ is a deterministic function of $S_{o}$, the data processing inequality [7, Ch. 2] gives
@@
Since the supports of $\{P_{a}\}_{a\in A}$ are pairwise disjoint, the chain rule for entropy [7, Ch. 2][8, Ch. 3] yields
@@
The output marginal is $P_{A}\bar{\pi}+P_{J}Q$, and concavity [7, Ch. 2][8, Ch. 3] of entropy gives
@@
Then $T$ is a deterministic function of $S_{o}$, the data processing inequality [7, Ch. 2] yields
@@
Step 3: achievability by homogenizing the redundant part. Let $P_{\hat{S}\mid A}^{*}$ achieve $R^{(A)}(D/P_{A};\,d_{\text{Cn}},\,\pi_{A})$ [7, Ch. 10] [8, Ch. 7], and let
@@
Since $T_{\delta}$ is a deterministic function of $S_{o}$, by the data processing inequality [7, Ch. 2], we have
@@
Since the supports of $\{P_{a}\}_{a\in A_{\delta}}$ are pairwise disjoint, the chain rule for entropy [7, Ch. 2][8, Ch. 3] gives
@@
A-D Proof of Theorem 5.2
@@
Since $T_{\delta}$ is a deterministic function of $S_{o}$, by the data processing inequality [7, Ch. 2], we have
$I(S_{o};\hat{S}_{o})\geq I(T_{\delta};\hat{S}_{o}).$
Following the same concavity argument as in the converse of Theorem 4.1 [7, Ch. 2][8, Ch. 1], we obtain
$I(T_{\delta};\hat{S}_{o})\geq P_{\delta}\,I(\mathsf{A}_{\delta,o};\hat{\mathsf{S}}_{o}),$
@@
For achievability, let $P_{\hat{S}\mid A_{\delta}}^{*}$ be an optimal test channel that achieves the rate-distortion function $R^{(A_{\delta})}(\cdot)$ on the right-hand side [7, Ch. 10] [8, Ch. 7], and extend it to all of $S_{O}$ by assigning to every $j\notin A_{\delta}$ the output marginal induced by $P_{\hat{S}\mid A_{\delta}}^{*}$. This keeps the distortion unchanged because the symbols outside $A_{\delta}$ are $\delta$-redundant, and the same entropy calculation as in Appendix A-B shows that the resulting mutual information is exactly
@@
A-E Proof of Theorem 6.1
All distortion functions used in the theorem are bounded single-letter distortions on finite alphabets. Therefore the standard source–channel separation theorem applies [7, Chs. 7,10] [8, Ch. 7].
For closure distortion $d_{\text{Cn}}$, the source-coding side is governed by $R_{\mathrm{sem}}(D;\,d_{\text{Cn}},\,P_{O})$[7, Ch. 10] [8, Ch. 7], and the channel-coding side is governed by the Shannon capacity C(W)C(W)[7, Ch. 7] [8, Ch. 6]. Hence a distortion level $D$ is achievable whenever
@@
Let $B:=\mathbf{1}[S_{o}\in A]$. Since $B$ is a deterministic function of $S_{o}$, by the chain rule for mutual information [7, Ch. 2][8, Ch. 3], we have
@@
Because the sets $R_{\text{Cn}}(a)$ are pairwise disjoint, $g$ is well-defined. By data processing [7, Ch. 2],
@@
Let $p_{A}:=\epsilon_{\text{Cn}}/P_{A}$. Applying the classical Fano inequality [7, Ch. 2][8, Ch. 3] to the $|A|$-ary source conditioned on $B=1$ gives